OpenStudy (anonymous):

~~~~~~~~FAN AND MEDAL!~~~~~~~~~~~~~~~~~~~ What are the exact solutions of x2 − 3x − 1 = 0?

OpenStudy (tkhunny):

Have you considered the Quadratic Formula? Maybe Completing the Square?

OpenStudy (anonymous):

I'm not good at it :(

OpenStudy (tkhunny):

Well, now is the time. It's really the same thing. Shall we use the formula or the process to build the formula? Your choice.

OpenStudy (anonymous):

Use the formula :)

OpenStudy (tkhunny):

Okay, it uses parameters a, b, and c. Can you identify those in the given equation?

OpenStudy (anonymous):

yeah wouldn't that be x^2 the 3x and 1

OpenStudy (tkhunny):

You have the right idea, but cut off too large of pieces. We need ONLY the coefficients. From the x^2 term, we get a = 1 From the x term, we get b = -3 From the constant term, we get c = -1 It's important to get the signs, and not just the numbers. So far, so good?

OpenStudy (anonymous):

Okay yup we're getting there

OpenStudy (tkhunny):

Okay, now your challenge for the evening. Let's see you write out the quadratic formula! Go!!!!!

OpenStudy (anonymous):

lolll idk the formula

OpenStudy (tkhunny):

Well, then why did you pick the method? That makes no sense. \(x = \dfrac{-b\pm\sqrt{b^{2} - 4ac}}{2a}\) I was going to help you with the substitution step, but since you made me write the formula, I await with eagerness your best substitution efforts.

OpenStudy (anonymous):

lol what? i didnt know what u meant when u were letting me choose but anyway